OpenStudy (anonymous):

when 36.0 grams of water is decomposed through electrolysis, 4.0 grams of hydrogen gas is produced. What mass of oxygen gas is also produced?

OpenStudy (anonymous):

Help!!

OpenStudy (anonymous):

Can you help @ParthKohl?

OpenStudy (abmon98):

\[H2O \rightarrow H2+1/2O2 \] \[Number of moles=Mass(g)/Molar Mass(g/mol) \] \[Number of moles of H2O=36.0/18.0=2.0\] \[Number of moles of H2=4.0/2.0=2.0\] Mole Ratio of Reactants to Products 1:1:1/2 2:2:x x-represent the number of moles of O2 cross multiplication to solve for x. x=2*1/2/1=1 mole Atomic weight of O:16 so if two atoms are present we would have 16*2=32g 1=Mass/32 g Mass=32

OpenStudy (anonymous):

Thank you so much!

OpenStudy (abmon98):

your most welcome :)

OpenStudy (jfraser):

for this kind of problem, you don't even need a reaction, if you understand what conservation of mass really means. 4g of hydrogen comes from 36g of water. If oxygen is the other product, it must weigh 32g. Conservation of mass must be satisfied